Electromagnetic radiation of wavelength 77.9 nm was projected on a gallium plate with a 6.86×10-19 J of work function. Calculate the kinetic energy of an emitted electron.
Answer #1

energy of the radiation (E) = hc/, h is planck's constant, c is speed of light and is wavelength

E = 6.625 * 10-34 * 3 * 108 / (77.9 * 10-9) = 2.55 * 10-18 J

We know that : K.E = E - Work function = 2.55 * 10-18 - 6.86 * 10-19 = (25.5 - 6.86) * 10-19 J = 1.864 * 10-18 J